OpenStudy (anonymous):

Planets in our solar system do not revolve around the sun in perfect circles. Their orbits are more like ovals that scientists describe as which of the following?

OpenStudy (anonymous):

HI... \[\Huge \color{red}{\ast~~\ast}~~\color{blue}{\ast~~\ast}~~\color{green}{\ast~~\ast}~~\color{orange}{\ast~~\ast}~~\color{purple}{\ast~~\ast}~~\color{salmon}{\ast~~\ast}~~\color{red}{\ast}\] \[\Huge \color{red}{\mathbb{WELCOME}}~~\color{orange}{\mathbb{TO}}~~\color{purple}{\mathbb{OPENSTUDY}}\]

OpenStudy (anonymous):

can you please list the "following"?? this is the maths section....so you might get a faster reply in the appropriate section... Look above and you would see, "Find more Subjects" and click on the subject you want.

OpenStudy (anonymous):

i meant click on the "Find more subjects" and a list pops up and click on the subject you want...